OpenStudy (anonymous):

Solve the equation log base 5 (4x+1)=log base 5 (2x+7)

jimthompson5910 (jim_thompson5910):

log base 5 (4x+1)=log base 5 (2x+7) 4x+1 = 2x + 7 4x+1-2x = 7 4x - 2x = 7 - 1 keep going to solve for x

jimthompson5910 (jim_thompson5910):

make sure you check any and all potential answers you get

OpenStudy (anonymous):

are both x the same?

jimthompson5910 (jim_thompson5910):

what do you mean

OpenStudy (anonymous):

am i going to get 1 answer or two?

jimthompson5910 (jim_thompson5910):

just one potential answer, but you should always check any and all potential answers you get

OpenStudy (anonymous):

im lost my you help?

jimthompson5910 (jim_thompson5910):

4x - 2x = 7 - 1 2x = 6 x = ???

OpenStudy (fibonaccichick666):

the phrase combine like terms may help you understand what @jim_thompson5910 is saying

OpenStudy (anonymous):

x=3

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

whats my next step?

OpenStudy (fibonaccichick666):

that's it

OpenStudy (anonymous):

okay may you explian this to me

jimthompson5910 (jim_thompson5910):

well you have the answer, so you're done, but you should always check your answers

jimthompson5910 (jim_thompson5910):

this is given log base 5 (4x+1)=log base 5 (2x+7)

jimthompson5910 (jim_thompson5910):

the logs are the same base, so the stuff inside the logs must be equal 4x+1 = 2x+7

jimthompson5910 (jim_thompson5910):

from here, we solve this as any other linear equation: get x all by itself on one side

OpenStudy (anonymous):

oh so you combined the like terms

jimthompson5910 (jim_thompson5910):

yes something like 2x and 5x are like terms if you add them, you get 2x+5x = 7x

OpenStudy (anonymous):

lol yess i get it now

OpenStudy (tkhunny):

"log base 5 (4x+1)=log base 5 (2x+7)" You check your answer for at least two reasons: 1) Errors can occur. 2) Fundamental things can go wrong. In this case, these two things are NOT QUITE equivalent. log base 5 (4x+1)=log base 5 (2x+7) and 4x+1 = 2x+7 This is a Domain problem. You can try any value you like in 4x+1 or 2x+7, but this is not so of the logarithms of those values. Try x = -1. 4(-1) + 1 = -4 + 1 = -3 -- Nothing wrong with that. log(-3) -- That's no good!